Find the inverse Laplace transform L−1{F(s)} of the given function.

F(s) = 6s^2−10s+16 / s(s^2+4)

Your answer should be a function of t. Enclose arguments of functions in parentheses. For example, sin(2x).

L^−1 {F(s)} = _______________-
1
Expert's answer
2020-07-28T17:56:30-0400

F(s) = (6s2-10s+16)/(s(s2+4))

by partial fraction

(6s2-10s+16)/(s(s2+4)) = A/s + (Bs+C)/(s2+4)

(6s2-10s+16) = A(s2+4)+(Bs+C)s

by comparing coefficients

A=4, B=2, C=-10

so F(s)= 4/s + (2s-10)/(s2+4)

= 4/s + 2s/(s2+4) - 10/(s2+4)

apply laplace inverse

f(t) = 4 + 2cos(2t) - 5sin(2t)

since laplace inverse(1/s)=1

laplace inverse(s/(s2+a2))=cos(at),

laplace inverse (a/(s2+a2))=sin(at).